# Hermite trace polynomials and chaos decompositions for the Hermitian Brownian motion

### v2

@@ 3.1. A function on the symmetric group
On the symmetric group $S_{0}(n)$, consider the function $\chi^{n+1}_{q}:\alpha\mapsto q^{\left|\alpha\right|}$, and extend it to a function on the group algebra $\mathbb{C}[S_{0}(n)]$. As is well-known (see, for example, [GGK13, KN21]), this function is positive semi-definite for
@@
and is not positive semi-definite for other values of $q$. It follows that these functions are positive for all $n$ if
@@
$\pi_{n,q}:\mathbb{C}[S_{0}(n)]\rightarrow\operatorname{\mathrm{End}}\left((\mathbb{C}^{N})^{\otimes(n+1)}\right),$
@@
It is well-known [GGK13] (see also [Bia01, Ker03]) that
@@
We now discuss the kernel of the normalized character $\chi_{q}$. See Section 4 of [Pro76] for a related discussion.
@@
    • Let $\pi\in\mathcal{PS}_{0}(n)$ be a partial permutation of $[0,n]$ with $N$ linear orbits. Denote $a_{0},\ldots,a_{N-1}$ and $b_{0},\ldots,b_{N-1}$ the initial, respectively, final elements of these orbits (recall that if an element does not belong to any actual orbit of $\pi$, we consider it as a single-element linear orbit, in which case $a_{i}=b_{i}$). Here $a_{0},b_{0}$ belong to the orbit containing 00. A permutation $\sigma\in S_{0}(N-1)$ naturally acts on the linear orbits of $\pi$ by concatenating them. Denote the result of this action by $\sigma\circ\pi\in S_{0}(n)$. Thus, $(\sigma\circ\pi)(b_{i})=a_{\sigma(i)}$, and $(\sigma\circ\pi)(x)=\pi(x)$ otherwise. Then
      $\mathcal{N}_{gr,1/N}=\mathrm{Span}\left(\sum_{\sigma\in S_{0}(N-1)}(-1)^{\left|\sigma\right|}\sigma\circ\pi:\pi\in\mathcal{PS}_{0}(n,N)\right)$
@@
      $\mathcal{N}_{gr,-1/N}=\mathrm{Span}\left(\sum_{\sigma\in S_{0}(N-1)}\sigma\circ\pi:\pi\in\mathcal{PS}_{0}(n,N)\right).$
@@
(a) and the equivalence between the first three entries in (b) are well-known. For the final entry, denote by $\bar{\pi}$ the permutation obtained by “closing” the orbits of $\pi$; that is, $\bar{\pi}(b_{i})=a_{i}$ and $\bar{\pi}(x)=\pi(x)$ otherwise. Also let $\alpha\in S(n)$ be defined by $\alpha(a_{i})=i$ and arbitrarily otherwise; thus, $\alpha$ maps $\left\{a_{0},\ldots,a_{N-1}\right\}$ bijectively onto $[0,N-1]$. Then a calculation shows that $\alpha^{-1}\sigma\alpha\bar{\pi}=\sigma\circ\pi$. Therefore
$\sum_{\sigma\in S(N-1)}(-1)^{\left|\sigma\right|}\sigma\circ\pi=\sum_{\sigma\in S(N-1)}(-1)^{\left|\sigma\right|}\alpha^{-1}\sigma\alpha\bar{\pi}\in\mathcal{N}_{gr,q}.$
The argument for the opposite inclusion is very close to the proof of Theorem 4.5 in [Pro76], in a somewhat different language. Start with $\alpha^{-1}\sigma_{n}^{(1/N)}\beta\alpha$ in the ideal. Possibly by replacing $\beta$ by its multiple by appropriate transpositions, we may assume that $10,1,\ldots,N-$ 1 lie in different cycles of $\beta$. Denote $a_{i}=\alpha^{-1}(i)$ for $10\leq i\leq N-$ 1. Then $a_{0},\ldots,a_{N-1}$ lie in different cycles of $\alpha^{-1}\beta\alpha$. So $\alpha^{-1}\beta\alpha=\bar{\pi}$, where linear orbits of $\pi$ are cycles of $\alpha^{-1}\beta\alpha$ with initial elements $a_{0},a_{1},\ldots,a_{N-1}$, and cyclic orbits of $\pi$ are the remaining cycles of $\alpha^{-1}\beta\alpha$. It follows that
@@ 3.2. Algebra structure
Let $\alpha\in S(n)$, $\beta\in S_{0}(k)$. Define $\sigma_{n,k}\in S(n+k)$ by
@@
Define $\alpha\cup\beta\in S_{0}(n+k)$ by
@@
That is, $\alpha\cup\beta$ is obtained by: combining the cycles of $\alpha$ and $\beta$ containing 00 into
@@
The subalgebra $\bigoplus_{n=0}^{\infty}\mathbb{C}[S(n)]$ is called the generic tensor algebra in [Rai14].
@@
    $\mathcal{N}_{gr,q}$ is an ideal for this multiplication. Consequently the multiplication factors through to the quotient
    $\mathcal{TP}_{q}=\bigoplus_{n=0}^{\infty}\mathbb{C}[S_{0}(n)]/\mathcal{N}_{gr,q}.$
@@
(a) and (b) are immediate. (c) follows from (1) since $\mathcal{N}_{gr,q}$ is an ideal for the usual group algebra multiplication. ∎
@@
For $\alpha\in S_{0}(n)$ and $S\subset[n]$, the restriction $\alpha|_{S^{c}}$ of a permutation is the permutation on $[0,n]\setminus S$ defined by $\alpha|_{S^{c}}(x)=\alpha^{m}(x)$, where $m=\min\left\{k>0\ |\ \alpha^{k}(x)\in S^{c}\right\}$.
@@
For $A,B\subseteq\mathbb{Z}$, $\left|A\right|=\left|B\right|$, denote $P^{A}_{B}$ the unique order-preserving bijection from $A$ to $B$, as well as (by abuse of notation) the corresponding bijection between the collections of permutations $S(A)$ and $S(B)$.
@@
More generally, for $\pi\in S(n)$ with the cycle structure $\pi\in\operatorname{\mathcal{P}}_{1,2}(n)$, define the contraction
@@
where $\ell=\left|\mathrm{Pair}\left(\pi\right)\right|=n-\left|\pi\right|$. Extend $C_{\pi}$ linearly to $\mathbb{C}[S_{0}(n)]$.
@@
$q^{\operatorname{\mathrm{cyc}}_{0}((\tau\alpha)|_{\mathrm{supp}\left(\tau\right)^{c}})-\operatorname{\mathrm{cyc}}_{0}(\tau\alpha)+1}=\begin{cases}q^{-1},&(ij)\text{ is a cycle in }\alpha,\\ 1,&(i),(j)\text{ are cycles in }\alpha,\\ 1,&i,j\text{ are consecutive elements}\\ &\text{in the same cycle of $\alpha$ of length at least }3,\\ q,&\text{otherwise.}\end{cases}$
@@
Let $q\in\mathcal{Z}\setminus\left\{0\right\}$. Each $C_{\pi}$ preserves $\mathcal{N}_{gr,q}$. Therefore it factors through to the quotient $\mathcal{TP}_{q}$.
@@
We consider the case $q=\frac{1}{N}$; for $q=-\frac{1}{N}$, the argument similar. We will use the representation in Proposition 3.1. Let $\eta\in\mathcal{N}_{gr,q}$. Since $\mathcal{N}_{gr,q}$ is an ideal, $\tau\eta\in\mathcal{N}_{gr,q}$. So it suffices to show that if $\pi\in\mathcal{PS}(n)$ and $\eta=\sum_{\sigma\in S(N)}(-1)^{\left|\sigma\right|}\sigma\circ\pi$, then for any $S\subset[0,n]$,
$\sum_{\sigma\in S(N)}q^{\operatorname{\mathrm{cyc}}_{0}((\sigma\circ\pi)|_{S^{c}})-\operatorname{\mathrm{cyc}}_{0}(\sigma\circ\pi)}(-1)^{\left|\sigma\right|}(\sigma\circ\pi)|_{S^{c}}\in\mathcal{N}_{gr,q}.$
Moreover, it suffices to take $S=\left\{s\right\}$. We consider two cases.
Suppose $s=a_{i}=b_{i}$. Then $(\sigma\circ\pi)|_{\left\{s\right\}^{c}}=(\sigma|_{\left\{i\right\}^{c}})\circ(\pi|_{\left\{s\right\}^{c}})$ and $\operatorname{\mathrm{cyc}}_{0}((\sigma\circ\pi)|_{\left\{s\right\}^{c}})=\operatorname{\mathrm{cyc}}_{0}(\sigma|_{\left\{i\right\}^{c}})$. Each $\sigma^{\prime}\in S(N-1)$ appears as $\sigma|_{\left\{i\right\}^{c}}$ $N$ times, once when $(i)$ is a cycle in $\sigma$ (so that $\sigma^{\prime}$ has one less cycle than $\sigma$), and $N-1$ times corresponding to $\sigma(i)=j$ for each $j\in[N]\setminus\left\{i\right\}$ (so that $\sigma^{\prime}$ has the same number of cycles as $\sigma$). Thus
$\begin{split}\sum_{\sigma\in S(N)}q^{\operatorname{\mathrm{cyc}}_{0}((\sigma\circ\pi)|_{S^{c}})-\operatorname{\mathrm{cyc}}_{0}(\sigma\circ\pi)}(-1)^{\left|\sigma\right|}(\sigma\circ\pi)|_{S^{c}}&=\sum_{\sigma^{\prime}\in S(N-1)}(N-1-q^{-1})(-1)^{\left|\sigma^{\prime}\right|}\sigma^{\prime}\circ(\pi|_{\left\{s\right\}^{c}})\\ &=-\sum_{\sigma^{\prime}\in S(N-1)}(-1)^{\left|\sigma^{\prime}\right|}\sigma^{\prime}\circ(\pi|_{\left\{s\right\}^{c}})\in\mathcal{N}_{gr,q}.\end{split}$
@@
and $\mathcal{L}$ the direct sum of these operators. Note that for $\pi\in\operatorname{\mathcal{P}}_{1,2}(n)$, $C_{\pi}$ is a product of several transposition-type contractions, and
$\mathcal{L}^{\ell}=\ell!\sum_{\begin{subarray}{c}\pi\in\operatorname{\mathcal{P}}_{1,2}(n)\\ \left|\pi\right|=n-\ell\end{subarray}}C_{\pi}.$
@@
$\mathcal{L}_{n,k}^{(\ell)}=\ell!\sum_{\begin{subarray}{c}\pi\in\operatorname{\mathcal{P}}_{1,2}(n,k)\\ \left|\pi\right|=n-\ell\end{subarray}}C_{\pi}.$
@@
For $\eta\in\bigoplus_{n=0}^{\infty}\mathbb{C}[S_{0}(n)]$, define the Wick product
$\mathnormal{W}\left(\eta\right)=\mathnormal{T}\left(e^{-\mathcal{L}}\eta\right)=\sum_{\ell=0}^{\infty}(-1)^{\ell}\frac{1}{\ell!}\mathnormal{T}\left(\mathcal{L}^{\ell}(\eta)\right)=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{n-\left|\pi\right|}\mathnormal{T}\left(C_{\pi}(\eta)\right).$
@@
(a) follows by composing (terminating) power series in $\mathcal{L}$. For (b), the argument is very similar to the standard one for Hermite polynomials [dSCV85]. We expand
$\begin{split}\mathnormal{W}\left(\alpha\right)\mathnormal{W}\left(\beta\right)&=\sum_{{\sigma_{1}}\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{n-\left|{\sigma_{1}}\right|}\mathnormal{T}\left(C_{\sigma_{1}}(\alpha)\right)\sum_{{\sigma_{2}}\in\operatorname{\mathcal{P}}_{1,2}(k)}(-1)^{k-\left|{\sigma_{2}}\right|}\mathnormal{T}\left(C_{\sigma_{2}}(\beta)\right)\\ &=\sum_{{\sigma_{1}}\in\operatorname{\mathcal{P}}_{1,2}(n)}\sum_{{\sigma_{2}}\in\operatorname{\mathcal{P}}_{1,2}(k)}(-1)^{n+k-\left|{\sigma_{1}}\right|-\left|{\sigma_{2}}\right|}\mathnormal{T}\left(C_{\sigma_{1}}(\alpha)\cup C_{\sigma_{2}}(\beta)\right)\\ &=\sum_{{\sigma_{1}}\in\operatorname{\mathcal{P}}_{1,2}(n)}\sum_{{\sigma_{2}}\in\operatorname{\mathcal{P}}_{1,2}(k)}(-1)^{n+k-\left|{\sigma_{1}}\right|-\left|{\sigma_{2}}\right|}\sum_{\tau\in\operatorname{\mathcal{P}}_{1,2}([\left|\mathrm{Sing}\left({\sigma_{1}}\right)\right|+\left|\mathrm{Sing}\left({\sigma_{2}}\right)\right|])}\mathnormal{W}\left(C_{\tau}(C_{\sigma_{1}}(\alpha)\cup C_{\sigma_{2}}(\beta))\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n+k)}\mathnormal{W}\left(C_{\pi}({\alpha\cup\beta})\right)\sum_{\begin{subarray}{c}S_{1}\subset\mathrm{Pair}\left(\pi|_{[n]}\right)\\ S_{2}\subset\mathrm{Pair}\left(\pi|_{[n+1,n+k]}\right)\end{subarray}}(-1)^{\left|S_{1}\right|+\left|S_{2}\right|}.\end{split}$

### v3

@@ 3.1. A function on the symmetric group
On the symmetric group $S_{0}(n)$, consider the function $\chi^{n+1}_{q}:\alpha\mapsto q^{\left|\alpha\right|}$, and extend it linearly to a function on the group algebra $\mathbb{C}[S_{0}(n)]$. As is well-known (see, for example, [16, 28]), this function is positive semi-definite for
@@
and is not positive semi-definite for other values of $q$. It follows that these functions are positive definite for all $n$ if
@@
$\pi_{n,1/N}:\mathbb{C}[S_{0}(n)]\rightarrow\operatorname{\mathrm{End}}\left((\mathbb{C}^{N})^{\otimes(n+1)}\right),$
@@
It is well-known [16] (see also [4, 25]) that
@@
We now discuss the kernel of the normalized character $\chi_{q}$. See Section 4 of [33] for a related discussion.
@@
    • Let $\pi\in\mathcal{PS}_{0}(n)$ be a partial permutation of $[0,n]$ with $N$ linear orbits. Denote $a_{0},\ldots,a_{N-1}$ and $b_{0},\ldots,b_{N-1}$ the initial, respectively, final elements of these orbits (recall that if an element does not belong to any actual orbit of $\pi$, we consider it as a single-element linear orbit, in which case $a_{i}=b_{i}$). Here $a_{0},b_{0}$ belong to the orbit containing 0. For a permutation $\sigma\in S_{0}(N-1)$, define $\sigma\circ\pi\in S_{0}(n)$ by
      $(\sigma\circ\pi)(x)=\begin{cases}a_{\sigma(i)},&x=b_{i},\\ \pi(x),&\text{otherwise}.\end{cases}$
      In other words, we use the cycles of $\sigma$ to concatenate the linear orbits of $\pi$. Then
      $\mathcal{N}_{gr,1/N,n}=\mathrm{Span}\left(\sum_{\sigma\in S_{0}(N-1)}(-1)^{\left|\sigma\right|}\sigma\circ\pi:\pi\in\mathcal{PS}_{0}(n,N)\right)$
@@
      $\mathcal{N}_{gr,-1/N,n}=\mathrm{Span}\left(\sum_{\sigma\in S_{0}(N-1)}\sigma\circ\pi:\pi\in\mathcal{PS}_{0}(n,N)\right).$
@@
(a) and the equivalence between the first three entries in (b) are well-known. For the equivalence between the final two entries, denote by $\bar{\pi}$ the permutation obtained by “closing” the orbits of $\pi$; that is, $\bar{\pi}(b_{i})=a_{i}$ and $\bar{\pi}(x)=\pi(x)$ otherwise. Also let $\alpha\in S(n)$ be defined by $\alpha(a_{i})=i$ and arbitrarily otherwise; thus, $\alpha$ maps $\left\{a_{0},\ldots,a_{N-1}\right\}$ bijectively onto [0,N1][0,N-1]. Then a calculation shows that $\alpha^{-1}\sigma\alpha\bar{\pi}=\sigma\circ\pi$. Therefore
$\sum_{\sigma\in S(N-1)}(-1)^{\left|\sigma\right|}\sigma\circ\pi=\sum_{\sigma\in S(N-1)}(-1)^{\left|\sigma\right|}\alpha^{-1}\sigma\alpha\bar{\pi}$
is in the ideal generated by $\sigma_{n}^{(1/N)}$. The argument for the opposite inclusion is very close to the proof of Theorem 4.5 in [33], in a somewhat different language. The ideal is spanned by the elements of the form $\alpha^{-1}\sigma_{n}^{(1/N)}\beta\alpha$ for $\alpha,\beta\in S_{0}(n)$. Possibly by replacing $\beta$ by its multiple by appropriate transpositions, we may assume that $0,1,\ldots,N-1$ lie in different cycles of $\beta$. Denote $a_{i}=\alpha^{-1}(i)$ for $0\leq i\leq N-1$. Then $a_{0},\ldots,a_{N-1}$ lie in different cycles of $\alpha^{-1}\beta\alpha$. So $\alpha^{-1}\beta\alpha=\bar{\pi}$, where linear orbits of $\pi$ are cycles of $\alpha^{-1}\beta\alpha$ with initial elements $a_{0},a_{1},\ldots,a_{N-1}$, and cyclic orbits of $\pi$ are the remaining cycles of $\alpha^{-1}\beta\alpha$. It follows that
@@ 3.2. Algebra structure
Let $\alpha\in S_{0}(n)$, $\beta\in S_{0}(k)$. Define $\sigma_{n,k}\in S(n+k)$ by
@@
Identify $\alpha,\beta$ with elements in $S_{0}(n+k)$ by letting them act on the first $(n+1)$, respectively $(k+1)$ elements. Define $\alpha\cup\beta\in S_{0}(n+k)$ by
@@
That is, $\alpha\cup\beta$ is obtained by: combining the cycles of $\alpha$ and $\beta$ containing 0 into a cycle
@@
The subalgebra $\bigoplus_{n=0}^{\infty}\mathbb{C}[S(n)]$ is called the generic tensor algebra in [34].
@@
    $\mathcal{N}_{gr,q}$ is an ideal for this multiplication. Consequently, for $q\in\mathcal{Z}$, the multiplication factors through to the quotient
    $\mathcal{TP}_{q}=\left(\bigoplus_{n=0}^{\infty}\mathbb{C}[S_{0}(n)]\right)/\mathcal{N}_{gr,q}\simeq\bigoplus_{n=0}^{\infty}\left(\mathbb{C}[S_{0}(n)]/\mathcal{N}_{gr,q,n}\right).$
@@
Associativity of multiplication follows from the description of the operation $\cup$ after equation (3). (b) is immediate. For (c), since $\mathcal{N}_{gr,q}$ is an ideal for the usual group algebra multiplication, it follows from (3) that it is also an ideal for the Definition 3.3 multiplication. The final expression follows from Proposition 3.1. ∎
@@
For $\alpha\in S_{0}(n)$ and $S\subset[n]$, the restriction $\alpha|_{S^{c}}$ of a permutation is the permutation on $0,n]\setminus S$ defined by $\alpha|_{S^{c}}(x)=\alpha^{m}(x)$, where $m=\min\left\{k>0\ |\ \alpha^{k}(x)\in S^{c}\right\}$.
@@
For $A,B\subseteq\mathbb{Z}$, $\left|A\right|=\left|B\right|$, denote $P^{A}_{B}$ the unique order-preserving bijection from $A$ to $B$. By abuse of notation, we will also denote $P^{A}_{B}$ the corresponding bijection between the collections of permutations $S(A)$ and $S(B)$ (it should really be $\sigma\mapsto(P^{A}_{B})\sigma(P^{A}_{B})^{-1}$).
@@
More generally, for $\pi\in\operatorname{\mathcal{P}}_{1,2}(n)$, define the contraction
@@
where $\ell=\left|\mathrm{Pair}\left(\pi\right)\right|=\left|\pi\right|$. Extend $C_{\pi}$ linearly to $\mathbb{C}[S_{0}(n)]$.
@@
[…]
@@
Let $q\in\mathcal{Z}\setminus\left\{0\right\}$, and $\tau\in S(n)$ a transposition. Each $C_{\tau}$ maps $\mathcal{N}_{gr,q,n}$ to $\mathcal{N}_{gr,q,n-2}$.
@@
We consider the case $q=\frac{1}{N}$; for $q=-\frac{1}{N}$, the argument is similar. We will use the representation in Proposition 3.1. Let $\eta\in\mathcal{N}_{gr,q,n}$. Since $\mathcal{N}_{gr,q,n}$ is an ideal, $\tau\eta\in\mathcal{N}_{gr,q,n}$. So it suffices to show that if $\pi\in\mathcal{PS}(n)$ and $\eta=\sum_{\sigma\in S(N)}(-1)^{\left|\sigma\right|}\sigma\circ\pi$, then for any $S\subset[0,n]$,
$\sum_{\sigma\in S(N)}q^{\operatorname{\mathrm{cyc}}_{0}((\sigma\circ\pi)|_{S^{c}})-\operatorname{\mathrm{cyc}}_{0}(\sigma\circ\pi)}(-1)^{\left|\sigma\right|}(\sigma\circ\pi)|_{S^{c}}\in\mathcal{N}_{gr,q,n-2}.$
By induction, it suffices to take $S$ to be a singleton, $S=\left\{s\right\}$. We consider two cases.
Suppose $s=a_{i}=b_{i}$. Then $(\sigma\circ\pi)|_{\left\{s\right\}^{c}}=(\sigma|_{\left\{i\right\}^{c}})\circ(\pi|_{\left\{s\right\}^{c}})$ and $\operatorname{\mathrm{cyc}}_{0}((\sigma\circ\pi)|_{\left\{s\right\}^{c}})=\operatorname{\mathrm{cyc}}_{0}(\sigma|_{\left\{i\right\}^{c}})$. Each $\sigma^{\prime}\in S(N-1)$ appears as $\sigma|_{\left\{i\right\}^{c}}$ NN times, once when $(i)$ is a cycle in $\sigma$ (so that $\sigma^{\prime}$ has one less cycle than $\sigma$), and N1N-1 times corresponding to $\sigma(i)=j$ for each $j\in[N]\setminus\left\{i\right\}$ (so that $\sigma^{\prime}$ has the same number of cycles as $\sigma$). Thus
$\begin{split}\sum_{\sigma\in S(N)}q^{\operatorname{\mathrm{cyc}}_{0}((\sigma\circ\pi)|_{S^{c}})-\operatorname{\mathrm{cyc}}_{0}(\sigma\circ\pi)}(-1)^{\left|\sigma\right|}(\sigma\circ\pi)|_{S^{c}}&=\sum_{\sigma^{\prime}\in S(N-1)}(N-1-q^{-1})(-1)^{\left|\sigma^{\prime}\right|}\sigma^{\prime}\circ(\pi|_{\left\{s\right\}^{c}})\\ &=-\sum_{\sigma^{\prime}\in S(N-1)}(-1)^{\left|\sigma^{\prime}\right|}\sigma^{\prime}\circ(\pi|_{\left\{s\right\}^{c}})\in\mathcal{N}_{gr,q,n-2}.\end{split}$
@@
is well defined as a map from $\mathbb{C}[S_{0}(n)]/\mathcal{N}_{gr,q,n}$ to $\mathbb{C}[S_{0}(n-2)]/\mathcal{N}_{gr,q,n-2}$. Denote by $\mathcal{L}$ the direct sum of these operators, which we may consider as a map from $\mathcal{TP}_{q}$ to itself. Note that for $\pi\in\operatorname{\mathcal{P}}_{1,2}(n)$, $C_{\pi}$ is a product of several transposition-type contractions, and
$\mathcal{L}^{\ell}_{n}=\ell!\sum_{\begin{subarray}{c}\pi\in\operatorname{\mathcal{P}}_{1,2}(n)\\ \left|\pi\right|=\ell\end{subarray}}C_{\pi}.$
@@
$\mathcal{L}_{n,k}^{(\ell)}=\ell!\sum_{\begin{subarray}{c}\pi\in\operatorname{\mathcal{P}}_{1,2}(n,k)\\ \left|\pi\right|=\ell\end{subarray}}C_{\pi}.$
@@
Let $\eta\in\bigoplus_{n=0}^{\infty}\mathbb{C}[S_{0}(n)]$. By analogy with equation 2 from Section 2.4, define the Wick product
$\mathnormal{W}\left(\eta\right)=\mathnormal{T}\left(e^{-\mathcal{L}}\eta\right)=\sum_{\ell=0}^{\infty}(-1)^{\ell}\frac{1}{\ell!}\mathnormal{T}\left(\mathcal{L}^{\ell}(\eta)\right)=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left|\pi\right|}\mathnormal{T}\left(C_{\pi}(\eta)\right).$
@@
(a) follows by composing (terminating) power series in $\mathcal{L}$. For (b), the argument is very similar to the standard one for Hermite polynomials [11]. We expand
$\begin{split}\mathnormal{W}\left(\alpha\right)\mathnormal{W}\left(\beta\right)&=\sum_{{\sigma_{1}}\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left|{\sigma_{1}}\right|}\mathnormal{T}\left(C_{\sigma_{1}}(\alpha)\right)\sum_{{\sigma_{2}}\in\operatorname{\mathcal{P}}_{1,2}(k)}(-1)^{\left|{\sigma_{2}}\right|}\mathnormal{T}\left(C_{\sigma_{2}}(\beta)\right)\\ &=\sum_{{\sigma_{1}}\in\operatorname{\mathcal{P}}_{1,2}(n)}\sum_{{\sigma_{2}}\in\operatorname{\mathcal{P}}_{1,2}(k)}(-1)^{\left|{\sigma_{1}}\right|+\left|{\sigma_{2}}\right|}\mathnormal{T}\left(C_{\sigma_{1}}(\alpha)\cup C_{\sigma_{2}}(\beta)\right)\\ &=\sum_{{\sigma_{1}}\in\operatorname{\mathcal{P}}_{1,2}(n)}\sum_{{\sigma_{2}}\in\operatorname{\mathcal{P}}_{1,2}(k)}(-1)^{\left|{\sigma_{1}}\right|+\left|{\sigma_{2}}\right|}\\ &\qquad\cdot\sum_{\tau\in\operatorname{\mathcal{P}}_{1,2}([\left|\mathrm{Sing}\left({\sigma_{1}}\right)\right|+\left|\mathrm{Sing}\left({\sigma_{2}}\right)\right|])}\mathnormal{W}\left(C_{\tau}(C_{\sigma_{1}}(\alpha)\cup C_{\sigma_{2}}(\beta))\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n+k)}\mathnormal{W}\left(C_{\pi}({\alpha\cup\beta})\right)\sum_{\begin{subarray}{c}S_{1}\subset\mathrm{Pair}\left(\pi|_{[n]}\right)\\ S_{2}\subset\mathrm{Pair}\left(\pi|_{[n+1,n+k]}\right)\end{subarray}}(-1)^{\left|S_{1}\right|+\left|S_{2}\right|}.\end{split}$